1)6 more than a number
2)5 apples fewer than Sally
3)The Quotient of 34 and a number
4)x+3x=15
5)x=-7
___
9
6)24=(-12)+ x-6

Answer by elizabeth2005(20)


1) six more than a number would be you adding six to something so therefore it would be...6+n
2) five apples fewer then sally woud be you having less apples then sally does so that means the equation would be 5-n
3) the quotion means that you are going to divide 34 by n
4) x+3x=15...
first of all you have to combine like terms
...there is not a number infront of x but it is understood to be one
so 1x+3x=4x...
now you have 4x=15..
you now need to divide both sides by 4
four divided by four cancells each other out and leaves you with x
and four divided by 15 leaves you with 3.75..
so x=3.75
now after you have found out what x is you need to check your work....
do it like this..say 3.75+3*3.75=15
5) x over nine = -7..
first of all you multiply each side by 9..
it leaves you with 9y=-63
you then divide by nine and it cancells out the nine...
so y = -63
6)...i am sorry but i dont clearly understand what you are wanting to know on this question im sorry